tyson:

o that would mean he did 21 hours that week. If he did 11 hours then I'd say he did 10% more, because 11 is 110% of 10, which would be 10% more.

tyson:

woops didnt mean to do that

tyson:

AngeI:

This is just 3/7 times 4

jhonyy9:

so bc. 4 × (1/4) = 4/4 = 1 from this result that you need just multiply the 3/7 by 4 and will get the right answer sure hope now this is understandably